OpenStudy (anonymous):

Inverse of f(x) = 3^x

OpenStudy (anonymous):

Consider: \[f(x)=y=3^x\] Replace the y with x and the x with y (since you are reflecting it about the line y=x. Giving: \[g(y)=x=3^y\] Take the log base 3 of both sides giving: \[\log_3(x)=\log_3(3^y)=y\] Using the identity: \[\log_a(a^b)=b\]